Sales are the function of advertising in The Dawn and Diva Magazine (X, Y). 

S = XY2

If the price of advertising in The Dawn and Diva Magazine is Rs.5 and Rs.10 respectively. The total budget for advertising is Rs.105. For maximizing the sales of Dawn and Diva Magazine find out the best combination of advertisement in newspapers and magazines by using Lagrangian multiplier



1
Expert's answer
2021-01-29T07:25:51-0500

The sales function is S = XY2

By a positive monotonic transformation, we get

lnS = lnX + 2lnY

Given:

Px = 5

Py= 10

M = 105

M ≥ PxX + PyY

105 ≥ 5X + 10 Y

Forming the Lagrangian, we get:

α = lnX + 2lnY + λ(105 – 5X -10Y)

λ is the Lagrangian multiplier

"\frac{∂α}{∂X} = \frac{1}{X}-5λ \\

5λX = 1 \\

\frac{∂α}{∂Y} = \frac{2}{Y}-10λ \\

5λY = 1 \\

\frac{∂α}{∂λ} = 105 - 5X -10Y = 0 \\

\frac{5X}{5Y}= 1 \\

X = Y \\

105 = 5X + 10Y"

X* = Y* = 7